OpenStudy (anonymous):

on definite integral [o,pi/4] evaluate d^tannx/cos^2(x) dx

OpenStudy (anonymous):

e^tanx

OpenStudy (anonymous):

u = tan x , du = sec^2 dx sec^2 is the same as 1 / cos ^2 dx so we have e^ u du , the integral is e^u , or e^(tanx) , now plug back in upper and lower limits. so we have e^(tan pi/4) - e^(tan 0) = e^1 - e^0 = e - 1

OpenStudy (anonymous):

integral e^ u = e^u + c , but we have upper and lower limits